# Is Mamba Effective for Time Series Forecasting?

### v2

@@ Introduction
Linear Tokenization Layer
The input for the Embedding layer is $U_{in}$. Similar to iTransformer (Liu et al. 2023), we commence by tokenizing the time series, a method analogous to the tokenization of sequential text in natural language processing, to standardize the temporal series format. This pivotal task is executed by a single linear layer (3).
$\bm{U}=\mathrm{Linear}(Batch(U_{in}))$ (3)
@@
Mamba VC Encoding Layer
Within this layer, our primary objective is to extract the VC by linking variates that exhibit analogous trends aiming to learn the mutual information therein, thereby enabling a more robust representation of variates. The Transformer architecture confers the capacity for global attention (Vaswani et al. 2017), enabling the computation of the impact of all other variables upon a given variable, which facilitates the learning of precise information. Nonetheless, with an escalation in the number of variables, the computational burden of this global attention—characterized by its quadratic complexity—becomes untenable. In contrast, Mamba’s selective mechanism can discern the significance of different variables akin to an attention mechanism, and it exhibits a computational overhead that escalates in a near-linear fashion with an increasing count of variables. Yet, the unilateral nature of Mamba precludes it from attending to global variables in the manner of the Transformer; its selection mechanism is unidirectional, capable only of incorporating antecedent variables. To surmount this limitation, we employ two Mamba blocks to be combined as a bidirectional Mamba layer (4), which facilitates the acquisition of correlations among all variables, both antecedent and subsequent.
$\overrightarrow{\bm{Y}}=\overrightarrow{\mathrm{Mamba~{}Block}}(\bm{U})$ (4)
$\overleftarrow{\bm{Y}}=\overleftarrow{\mathrm{Mamba~{}Block}}(\bm{U})$
@@
FFN TD Encoding Layer
At this layer, we further process the output of the Mamba VC Encoding Layer. Firstly, we employ a normalization layer (Liu et al. 2023) to enhance convergence and training stability in deep networks by standardizing all variates to a Gaussian distribution, thereby minimizing disparities resulting from inconsistent measurements. Then, the feed-forward network (FFN) is used on the series representation of each variate. The FFN layer encodes observed time series and decodes future series representations using dense non-linear connections. During this procedure, FFN implicitly encodes TD. Finally, another normalization layer is set to adjust the future series representations as line 11-13 in Alg.1.
Projection Layer
@@
Experiments
Datasets and Baselines
We conduct extensive experiments on several real-world datasets, including Electricity, Exchange-Rate (Lai et al. 2018), Traffic, Weather, Solar-Energy, PEMS (PEMS03, PEMS04, PEMS07, PEMS08), ETT (Electricity Transformer Temperature) (Zhou et al. 2021) (ETTh1, ETTh2, ETTm1, ETTm2). All of them are multivariate time series datasets adopted by Liu et al. (2023). Tab. 1 shows the statistics of these datasets.
Our models are fairly compared with 6 representative and state-of-the-art forecasting models, including (1) Transformer-based methods: Autoformer (Wu et al. 2021), FEDformer (Zhou et al. 2022), PatchTST (Nie et al. 2022), and (2) Linear-based methods: DLinear (Zeng et al. 2023), RLinear (Li et al. 2023b).
Table 1: The statistics of the datasets.
@@
Table 2: Full results of S-Mamba and balines. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 12, 24, 48, 96 for PEMS and 96, 192, 336, 720 for other datasets. Results are averaged from all forecast lengths and full Results are in the Appendix A. The best results are highlighted in bold red font, while the second best results are presented in underlined blue font. The results of baselines are reported by Liu et al. (2023).
@@
ECL 0.170 0.265 0.178 0.270 0.219 0.298 0.205 0.290 0.244 0.334 0.251 0.344 0.192 0.295 0.212 0.300 0.214 0.327 0.227 0.338
Exchange 0.367 0.408 0.360 0.403 0.378 0.417 0.367 0.404 0.940 0.707 0.370 0.413 0.416 0.443 0.354 0.414 0.519 0.429 0.613 0.539
Traffic 0.414 0.276 0.428 0.282 0.626 0.378 0.481 0.304 0.550 0.304 0.760 0.473 0.620 0.336 0.625 0.383 0.610 0.376 0.628 0.379
Weather 0.251 0.276 0.258 0.278 0.272 0.291 0.259 0.281 0.259 0.315 0.271 0.320 0.259 0.287 0.265 0.317 0.309 0.360 0.338 0.382
Solar-Energy 0.240 0.273 0.233 0.262 0.369 0.356 0.270 0.307 0.641 0.639 0.347 0.417 0.301 0.319 0.330 0.401 0.291 0.381 0.885 0.711
PEMS03 0.122 0.228 0.113 0.221 0.495 0.472 0.180 0.291 0.169 0.281 0.326 0.419 0.147 0.248 0.278 0.375 0.213 0.327 0.667 0.601
PEMS04 0.103 0.211 0.111 0.221 0.526 0.491 0.195 0.307 0.209 0.314 0.353 0.437 0.129 0.241 0.295 0.388 0.231 0.337 0.610 0.590
PEMS07 0.089 0.188 0.101 0.204 0.504 0.478 0.211 0.303 0.235 0.315 0.380 0.440 0.124 0.225 0.329 0.395 0.165 0.283 0.367 0.451
PEMS08 0.148 0.224 0.150 0.226 0.529 0.487 0.280 0.321 0.268 0.307 0.441 0.464 0.193 0.271 0.379 0.416 0.286 0.358 0.814 0.659
ETTm1 0.398 0.405 0.407 0.410 0.414 0.407 0.387 0.400 0.513 0.496 0.419 0.419 0.400 0.406 0.403 0.407 0.448 0.452 0.588 0.517
ETTm2 0.288 0.332 0.288 0.332 0.286 0.327 0.281 0.326 0.757 0.610 0.358 0.404 0.291 0.333 0.350 0.401 0.305 0.349 0.327 0.371
ETTh1 0.455 0.450 0.454 0.447 0.446 0.434 0.469 0.454 0.529 0.522 0.541 0.507 0.458 0.450 0.456 0.452 0.440 0.460 0.496 0.487
ETTh2 0.381 0.405 0.383 0.407 0.374 0.398 0.387 0.407 0.942 0.684 0.611 0.550 0.414 0.427 0.559 0.515 0.437 0.449 0.450 0.459
Overall Performance
Tab.2 presents a comparative analysis of the overall performance of our models and other baseline models across all datasets. From the data presented in the table, we summarize two observations and attach the analysis: (1) S-Mamba has attained commendable outcomes on the Electricity, Traffic, PEMS, PEMS and Solar datasets. These datasets are distinguished by their numerous variates, most of which have the same period. It is worth noting that variates with the same period are more likely to contain learnable VC. Mamba VC Fusion Layer benefits from this characteristic and improves S-Mamba performance. (2) In the context of the ETT, and Exchange data sets, S-Mamba did not demonstrate a pronounced superiority in performance; indeed, it exhibited a suboptimal outcome. This can be attributed to the fact that these datasets are characterized by a limited number of variates, predominantly of an aperiodic nature. Consequently, there are weak VCs between these variates, and the employment of VC by S-Mamba may inadvertently introduce noise into the predictive model, thus impeding its predictive accuracy. (3) The Weather data set is special in that it has fewer variables and most variables are aperiodic, but S-Mamba still achieves the best performance on it.
Refer to captionRefer to captionRefer to captionRefer to caption
(a) S-Mamba
Refer to captionRefer to captionRefer to captionRefer to caption
(b) iTransformer
Figure 3: Comparison of forecasts between S-Mamba and iTransformer on four datasets when the input length is 96 and the forecast length is 96. The blue line represents the ground truth and the red line represents the forecast.
Furthermore, to provide a more intuitive assessment of S-Mamba’s forecast capabilities, we visually compare the predictions of S-Mamba and the leading baseline, iTransformer, on four datasets: Electricity, Weather, Traffic, and Exchange, through graphical representation. Specifically, we randomly select a variate and then input its lookback sequence, where the true subsequent sequence is depicted as a blue line and the model’s forecast is represented by a red line in Fig.3. It is evident that on the ECL, Weather, and Traffic datasets, S-Mamba’s predictions closely approximate the actual values, with nearly perfect alignment observed on the ECL and Traffic datasets. However, on the Exchange dataset, both models exhibit relatively modest performance.
(a) Electricity
(b) Traffic
Figure 4: Comparison of S-Mamba and five baselines on MSE, Training Time, and GPU Memory.
Model Efficiency
To evaluate the computational efficiency of the models, We compare the memory usage and computing time of S-Mamba with several baselines on the ECL and Traffic. Independent runs are conducted on a single NVIDIA RTX3090 GPU and meticulously document the results in Fig.4. In our analysis, a bubble chart to depict the measurement outcomes, wherein the vertical axis denotes the Mean Squared Error (MSE), the horizontal axis quantifies the training duration, and the bubble magnitude correlates with the video memory utilization. The visualization reveals that the S-Mamba algorithm attains the most favorable MSE metric across both the Electricity and Traffic datasets. When benchmarked against Transformer-based models, S-Mamba typically necessitates a reduced training time while concurrently minimizing GPU memory consumption. While the RLinear model does utilize minimal GPU memory and curtails training time, it does not confer a competitive edge in terms of predictive precision. Overall, S-Mamba manifests exemplary predictive accuracy with a low computational resource footprint.
Table 3: Ablation study on ECL, Traffic, Weather, and Solar-Energy. The lookback length $S=96$, while the forecast length $T\in\left\{96,192,336,720\right\}$. Results are averaged from all forecast lengths.
Design Variate Temporal ECL Traffic Weather Solar-Energy
MSE MAE MSE MAE MSE MAE MSE MAE
S-Mamba Mamba FFN 0.170 0.265 0.411 0.273 0.251 0.276 0.240 0.273
Replace Mamba uni-Mamba 0.182 0.280 0.538 0.342 0.244 0.268 0.250 0.282
Mamba bi-Mamba 0.182 0.281 0.551 0.371 0.246 0.271 0.255 0.282
Mamba Attention 0.179 0.278 0.561 0.379 0.247 0.272 0.261 0.288
Attention FFN 0.178 0.270 0.428 0.282 0.258 0.279 0.233 0.262
w/o Mamba w/o 0.174 0.269 0.418 0.278 0.253 0.279 0.247 0.278
w/o FFN 0.193 0.276 0.461 0.294 0.265 0.283 0.261 0.283
Ablation Study
To evaluate the efficacy of the components within S-Mamba, we conduct ablation studies by substituting or eliminating the VC and TD encoding layers. Specifically, the VC encoding layer was replaced with an Attention mechanism or entirely removed. This modification is predicated on the empirical evidence from iTransformer experiments (Liu et al. 2023), which demonstrate that Attention was the optimal encoder for VC. Conversely, the TD encoding layer is replaced with Attention, bidirectional Mamba, unidirectional Mamba, or omitted altogether. The choice of bidirectional Mamba, which is set to benchmark Attention, is made to facilitate global temporal information extraction. The rationale behind employing unidirectional Mamba is its resemblance to RNN models, so inherently possesses the capacity to preserve sequential relationships, thereby making it a suitable candidate for evaluating the impact of sequential encoding on TD.
Our experimental investigations are conducted on four datasets: Electricity, Traffic, Weather, and Solar Energy. The findings from these experiments in Tab.6 indicate that Mamba exhibite superior performance in VC encoding, whereas the Feed-Forward Network (FFN) maintained its dominance in TD encoding.
Can Variate Order Affect Mamba’s Performance?
In the context of the S-Mamba model, the existence of the Hippo matrix (Gu et al. 2020) theoretically enables the retention of information over an indefinite length. However, this very feature predisposes the model to initially prioritize proximal information. S-Mamba employs a bidirectional Mamba layer to extract VC. This initial bias towards neighboring variates could potentially impede the acquisition of a global VC, especially when relevant variates are distant from each other. This consideration prompted us to investigate the impact of variate ordering on the performance of S-Mamba.
We hypothesize that within the dataset, variates are more likely to extract valuable information from other periodic variates. Conversely, the propensity to acquire noise increases when interacting with aperiodic variates. To address this, we apply the Fourier transform (Bracewell 1989) to help categorize variates into periodic and aperiodic groups. Our experimentation focuses on the ECL dataset, notable for its numerous periodic and aperiodic variates. The initial arrangement of these variates is depicted in the accompanying Fig.5.
Subsequent trials involve repositioning the aperiodic variates towards the middle or end of the variates sequence, followed by evaluating the predictive capabilities of the models trained on these modified datasets. For comparative analysis, we also included experiments with the iTransformer. The variate distribution and corresponding model performance are illustrated in the Fig. . Our findings suggest that the S-Mamba model’s performance remains largely unaffected by the perturbation of variate order. This implies that through adequate training, the S-Mamba can effectively mitigate the initial limitation of the Hippo matrix, which biases attention towards nearby information.
Table 4: Performance comparison between Autoformer, Flashformer and Flowformer and their variants with Mamba encoder block on ECL and Traffic datasets, where we fixed the lookback length $L=96$ and forecasting length $T=96$.
Datasets ECL Traffic
Models Auto Auto-M Flash Flash-M Flow Flow-M Auto Auto-M Flash Flash-M Flow Flow-M
Memory(GiB) 1.59 1.22 0.91 0.63 0.83 0.68 1.68 1.35 1.16 0.80 1.06 0.79
Speed(ms/iters) 54.6 45.1 62.8 40.6 57.4 38.0 126.1 97.3 110.3 91.5 101.5 81.2
MSE 0.201 0.191 0.253 0.252 0.252 0.246 0.613 0.607 0.68 0.67 0.658 0.649
MAE 0.317 0.305 0.347 0.345 0.349 0.344 0.338 0.381 0.374 0.465 0.360 0.365
Can Mamba Outperform Advanced Transformers?
Beyond the foundational Transformer architecture, some advanced Transformers have been introduced, predominantly focusing on the augmentation of the self-attention mechanism. We aim to determine whether Mamba can still maintain an advantage over these advanced Transformers. To this end, we conduct a comparative experiment in which we directly replace the Encoder layer of three advanced transformers: Autoformer (Wu et al. 2021), Flashformer (Dao et al. 2022) and Flowformer (Wu et al. 2022) with a unidirectional Mamba for TSF tasks to get Trans-M, Auto-M, Fed-M, Flash-M and Flow-M and and compare their performance. The results are astonishing as Tab.4, which straightforward modification with Mamba enhances the performance of these models and also reduces GPU memory usage and running time.
Figure 5: In the ECL dataset, adjust the distribution of periodic and aperiodic variates. The left side represents the distribution, while the right side indicates models’ performance.
Can S-Mamba Benefit from Increasing Lookback Length?
Prior research has shown that the performance of Transformer-based models does not consistently improve with increasing lookback sequence length $L$, which is somewhat unexpected. A plausible explanation is that the temporal sequence relationship is overlooked under the self-attention mechanism, as it disregards the sequential order, and in some instances, even inverts it. Mamba, resembling a Recurrent Neural Network (Medsker, Jain et al. 2001), concentrates on the preceding window during information extraction, thereby preserving certain sequential attributes. This prompts an exploration of Mamba’s potential effectiveness in temporal sequence information fusion, aiming to address the issue of diminishing or stagnant performance with increasing lookback length. Consequently, we add an additional Mamba block between the encoder Layer and decoder layer of Transformers. The role of this Mamba Block is to add a layer of time sequence dependence from the information output by the encoder layer, to add some information similar to position embedding before the decoder layer processes it. We experiment with Autoformer (Wu et al. 2021), Transformer (Vaswani et al. 2017), and Informer (Zhou et al. 2021) to get Auto-M, Infor-M, and Refor-M, and evaluate their performance with varying lookback lengths. We also test the performance of S-Mamba as the lookback length changes. The results are in the Fig.6, from which we can observe three results. (1) S-Mamba can enhance its performance as the input lengthens, but we believe this is not solely due to the Mamba Block, but rather to the Variate-based Embedding method proposed by iTransformer (Liu et al. 2023). (2) With the addition of a Mamba Block, all three models show performance improvements. Notably, the performance of Informer has seen the most significant enhancement. (3) Despite these variants’ performance gains sometimes, they do not achieve optimization with longer lookback lengths. This is consistent with the findings of Zeng et al. (2023), which also suggest that encoding temporal sequence information into the model beforehand does not resolve this issue.
Figure 6: Forecasting performance on ECL and Traffic with the lookback length $L\in\{96,192,336,720\}$ while the prediction length $S=96$.
Is Mamba Generalizable in TSF?
The emergence of pre-trained models (Devlin et al. 2018) and large language models (Chang et al. 2023) based on the Transformer architecture has underscored the Transformer’s ability to discern similar patterns across diverse data, highlighting its generalization capabilities. In the context of TSF, it is observed that all variates exhibit a similar pattern of differences, for example, temperature and humidity in weather forecasts always have similar trends. In other words, the generalization potential of the Transformer for sequence data may also take effect on TSF tasks.
In this vein, iTransformer (Liu et al. 2023) conducts a pivotal experiment. The study involves masking a majority of the variates in a dataset and training the model on a limited subset of variates. Subsequently, the model was tasked with forecasting all variates, including those previously unseen, based on the learned information from the few variates. Building on this, we seek to evaluate the generalization capabilities of Mamba in TSF tasks. An experiment is proposed wherein the S-Mamba are trained on merely 40% of the variates in the PEMS03, PEMS04, ECL, Weather, Traffic, and Solar datasets. Then they are employed to predict 100% variates, and the results are subjected to statistical analysis. The outcomes of this investigation in 7 reveal that S-Mamba exhibits generalization potential in the six datasets, which proves their generalizability in TSF tasks.
Conlusion
Transformer-based models have consistently exhibited outstanding performance in the field of time series forecasting (TSF), while Mamba has recently gained popularity, and has been shown to surpass the Transformer in various domains by delivering superior performance while reducing memory and computational overhead. Motivated by these advancements, we sought to investigate the potential of Mamba-based models in the TSF domain, with the aim of uncovering new research avenues for this field. To this end, we introduced a Mamba-based model, Simple-Mamba (S-Mamba), which transfers the task of variates information fusion from the Transformer architecture to the Mamba block. We evaluated their performance across a multitude of datasets, and our findings indicate that S-Mamba not only reduces GPU memory and computational overhead but also achieves leading performance. Furthermore, we conduct extensive experiments to prove Mamba possesses robust capabilities and exhibits remarkable potential to replace the Transformer in the TSF tasks.
Appendix A Full Results
We compare the representative and the state-of-the-art models with S-Mamba under different prediction lengths in Tab.5. The lookback length $L$ is set to 96 for all models. The number on the left side of the table represents the length of the output sequence $T$. The optimal results are highlighted in bold red font, while the suboptimal results are presented in underlined blue font. The results of baselines are reported by Liu et al. (2023).
The full results of the ablation study are represented in Tab.6.
Table 5: Full results of the long-term time series forecasting task.
@@ Appendix A Full Results
ECL 96 0.139 0.235 0.148 0.240 0.201 0.281 0.181 0.270 0.219 0.314 0.237 0.329 0.168 0.272 0.197 0.282 0.193 0.308 0.201 0.317
@@
Table 6: Full results of ablation study.

### v3

@@ 4 Methodology
4.1 Linear Tokenization Layer
The input for the Linear Tokenization Layer is $U_{in}$. Similar to iTransformer [37], we commence by tokenizing the time series, a method analogous to the tokenization of sequential text in natural language processing, to standardize the temporal series format. This pivotal task is executed by a single linear layer in Eq. (3).
$\bm{U}=\mathrm{Linear}(Batch(U_{in})),$ (3)
@@ 4.1 Linear Tokenization Layer
4.2 Mamba VC Encoding Layer
Within this layer, the primary objective is to extract the VC by linking variates that exhibit analogous trends aiming to learn the mutual information therein. The Transformer architecture confers the capacity for global attention [52], enabling the computation of the impact of all other variates upon a given variate, which facilitates the learning of precise information. However, the computational load of global attention escalates exponentially with an increase in the number of variates, potentially rendering it impractical. This limitation could restrict the application of Transformer-based algorithms in real-world scenarios. In contrast, Mamba’s selective mechanism can discern the significance of different variates akin to an attention mechanism, and it exhibits a computational overhead that escalates in a near-linear fashion with an increasing count of variates. Yet, the unilateral nature of Mamba precludes it from attending to global variates in the manner of the Transformer; its selection mechanism is unidirectional, capable only of incorporating antecedent variates. To surmount this limitation, we employ two Mamba blocks to be combined as a bidirectional Mamba layer as Eq. (4), which facilitates the acquisition of correlations among all variates.
$\overrightarrow{\bm{Y}}=\overrightarrow{\mathrm{Mamba~{}Block}}(\bm{U}),$ (4)
$\overleftarrow{\bm{Y}}=\overleftarrow{\mathrm{Mamba~{}Block}}(\bm{U}).$
@@ 4.2 Mamba VC Encoding Layer
4.3 FFN TD Encoding Layer
At this layer, we further process the output of the Mamba VC Encoding Layer. Firstly, we employ a normalization layer [37] to enhance convergence and training stability in deep networks by standardizing all variates to a Gaussian distribution, thereby minimizing disparities resulting from inconsistent measurements. Then, the feed-forward network (FFN) is used on the series representation of each variate. The FFN layer encodes observed time series and decodes future series representations using dense non-linear connections. During this procedure, FFN implicitly encodes TD by keeping the sequential relationships. Finally, another normalization layer is set to adjust the future series representations.
4.4 Projection Layer
@@
5 Experiments
5.1 Datasets and Baselines
We conduct experiments on thirteen real-world datasets. For convenience of comparison, we divide them into three types. (1) Traffic-related datasets: Traffic [59] and PEMS [10]. Traffic is a collection of hourly road occupancy rates from the California Department of Transportation, capturing data from 862 sensors across San Francisco Bay area freeways from January 2015 to December 2016. And PEMS is a complicated spatial-temporal time series for public traffic networks in California including four public subsets (PEMS03, PEMS04, PEMS07, PEMS08), which are the same as SCINet [36]. Traffic-related datasets are characterized by a large number of variates, most of which are periodic. (2) ETT datasets: ETT [71] (Electricity Transformer Temperature) comprises data on load and oil temperature, collected from electricity transformers over the period from July 2016 to July 2018. It contains four subsets, ETTm1, ETTm2, ETTh1 and ETTh2. ETT datasets have few varieties and weak regularity. (3) Other datasets: Electricity [59], Exchange [59], Weather [59], and Solar-Energy [29]. Electricity records the hourly electricity consumption of 321 customers from 2012 to 2014. Exchange collects daily exchange rates of eight countries from 1990 to 2016. Weather contains 21 meteorological indicators collected every 10 minutes from the Weather Station of the Max Planck Biogeochemistry Institute in 2020. Solar-Energy contains solar power records in 2006 from 137 PV plants in Alabama State which are sampled every 10 minutes. Among them, the Electricity and Solar-Energy datasets contain many variates, and most of them are periodic, while the Exchange and Weather datasets contain fewer variates, and most of them are aperiodic. Tab. 1 shows the statistics of these datasets.
Table 1: The statistics of the thirteen public datasets.
@@ 5.1 Datasets and Baselines
Electricity 321 26,304 1hour
Exchange 8 7,588 1day
Weather 21 52,696 10min
Solar-Energy 137 52,560 10min
Table 2: Full results of S-Mamba and baselines on traffic-related datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 12, 24, 48, 96 for PEMS and 96, 192, 336, 720 for Traffic.
@@
Avg 0.414 0.276 0.428 0.282 0.626 0.378 0.481 0.304 0.550 0.304 0.760 0.473 0.620 0.336 0.625 0.383 0.610 0.376 0.628 0.379
Avg 0.122 0.228 0.113 0.221 0.495 0.472 0.180 0.291 0.169 0.281 0.326 0.419 0.147 0.248 0.278 0.375 0.213 0.327 0.667 0.601
Avg 0.103 0.211 0.111 0.221 0.526 0.491 0.195 0.307 0.209 0.314 0.353 0.437 0.129 0.241 0.295 0.388 0.231 0.337 0.610 0.590
Avg 0.089 0.188 0.101 0.204 0.504 0.478 0.211 0.303 0.235 0.315 0.380 0.440 0.124 0.225 0.329 0.395 0.165 0.283 0.367 0.451
Avg 0.148 0.224 0.150 0.226 0.529 0.487 0.280 0.321 0.268 0.307 0.441 0.464 0.193 0.271 0.379 0.416 0.286 0.358 0.814 0.659
Our models are fairly compared with 9 representative and state-of-the-art (SOTA) forecasting models, including (1) Transformer-based methods: iTransformer [37], PatchTST [44], Crossformer [69], FEDformer [72], Autoformer [59]; (2) Linear-based methods: RLinear [33], TiDE [14], DLinear [66]; and (3) Temporal Convolutional Network-based methods: TimesNet [57]. The brief introductions of these models are as follows:
  • iTransformer reverses the order of information processing, which first analyzes the time series information of each individual variate and then fuses the information of all variates. This unique approach has positioned iTransformer as the current SOTA model in TSF.
  • PatchTST segments time series into subseries patches as input tokens and uses channel-independent shared embeddings and weights for efficient representation learning.
  • Crossformer introduces a cross-attention mechanism that allows the model to interact with information between different time steps to help the model capture long-term dependencies in time series.
  • FEDformer is a frequency-enhanced Transformer that takes advantage of the fact that most time series tend to have a sparse representation in well-known basis such as Fourier transform to improve performance.
  • Autoformer takes a decomposition architecture that incorporates an auto-correlation mechanism and updates traditional sequence decomposition into the basic inner blocks of the depth model.
  • RLinear is the SOTA linear model, which employs reversible normalization and channel independence into pure linear structure.
  • TiDE is a Multi-layer Perceptron (MLP) based encoder-decoder model.
  • DLinear is the first linear model in TSF and a simple one-layer linear model with decomposition architecture.
  • TimesNet uses TimesBlock as a task-general backbone, transforms 1D time series into 2D tensors, and captures intraperiod and interperiod variations using 2D kernels.
Table 3: Full results of S-Mamba and baselines on ETT datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720.
@@
Avg 0.398 0.405 0.407 0.410 0.414 0.407 0.387 0.400 0.513 0.496 0.419 0.419 0.400 0.406 0.403 0.407 0.448 0.452 0.588 0.517
Avg 0.288 0.332 0.288 0.332 0.286 0.327 0.281 0.326 0.757 0.610 0.358 0.404 0.291 0.333 0.350 0.401 0.305 0.349 0.327 0.371
Avg 0.455 0.450 0.454 0.447 0.446 0.434 0.469 0.454 0.529 0.522 0.541 0.507 0.458 0.450 0.456 0.452 0.440 0.460 0.496 0.487
Avg 0.381 0.405 0.383 0.407 0.374 0.398 0.387 0.407 0.942 0.684 0.611 0.550 0.414 0.427 0.559 0.515 0.437 0.449 0.450 0.459
Table 4: Full results of S-Mamba and baselines on Electricity, Exchange, Weather and Solar-Energy datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720.
Electricity 96 0.139 0.235 0.148 0.240 0.201 0.281 0.181 0.270 0.219 0.314 0.237 0.329 0.168 0.272 0.197 0.282 0.193 0.308 0.201 0.317
@@
Avg 0.240 0.273 0.233 0.262 0.369 0.356 0.270 0.307 0.641 0.639 0.347 0.417 0.301 0.319 0.330 0.401 0.291 0.381 0.885 0.711
5.2 Overall Performance
Tab. 2, Tab. 3, and Tab. 4 present a comparative analysis of the overall performance of our models and other baseline models across all datasets. The best results are highlighted in bold red font, while the second best results are presented in underlined purple font. From the data presented in these tables, we summarize three observations and attach the analysis: (1) S-Mamba attain commendable outcomes on the traffic-related, Electricity, and Solar-Energy datasets. These datasets are distinguished by their numerous variates, most of which are periodic. It is worth noting that period variates are more likely to contain learnable VC. Mamba VC Fusion Layer benefits from this characteristic and improves S-Mamba performance. (2) In the context of the ETT, and Exchange datasets, S-Mamba does not demonstrate a pronounced superiority in performance; indeed, it exhibits a suboptimal outcome. This can be attributed to the fact that these datasets are characterized by a few number of variates, predominantly of an aperiodic nature. Consequently, there are weak VCs between these variates, and the employment of Mamba VC Encoding layer by S-Mamba can’t bring useful information and even may inadvertently introduce noise into the predictive model, thus impeding its predictive accuracy. (3) The Weather dataset is special in that it has fewer variates and most variates are aperiodic, but S-Mamba still achieves the best performance on it. We think that this phenomenon arises from the tendency of variates in the Weather dataset to exhibit simultaneous trends of either falling or rising despite the absence of periodic patterns So the Mamba VC Encoding Layer of S-Mamba can still benefit from these data. Moreover, the Weather dataset exhibits large sections of rising or falling trends. The Feed-Forward Network (FFN) layer accurately records these relationships, which is also beneficial for S-Mamba’s comprehension.
(a) S-Mamba
(b) iTransformer
Figure 4: Comparison of forecasts between S-Mamba and iTransformer on five datasets when the input length is 96 and the forecast length is 96. The blue line represents the ground truth and the red line represents the forecast.
Furthermore, to provide a more intuitive assessment of S-Mamba’s forecast capabilities, we visually compare the predictions of S-Mamba and the leading baseline, iTransformer, on four datasets: Electricity, Weather, Traffic, Exchange, and ETTh1, through graphical representation. Specifically, we randomly select a variate and then input its lookback sequence, where the true subsequent sequence is depicted as a blue line and the model’s forecast is represented by a red line in Fig. 4. It is evident that on the Electricity, Weather, and Traffic datasets, S-Mamba’s predictions closely approximate the actual values, with nearly perfect alignment observed on the Electricity and Traffic datasets and are better than iTransformer. On the Exchange and ETTh1, the two models exhibit similar performance because the two datasets contain few variates, so there is no evident gap between using bidirectional Mamba or using Transformer for information fusion between variates.
Figure 5: Comparison of S-Mamba and six baselines on MSE, Training Time, and GPU Memory. The lookback length $L=96$, and the forecast length $T=12$ for PEMS07 and $T=96$ for other datasets
5.3 Model Efficiency
To evaluate the computational efficiency of the models, we compare the memory usage and computing time of S-Mamba with several baselines on PEMS07, Electricity, Traffic, and ETTm1. Independent runs are conducted on a single NVIDIA RTX3090 GPU with the batch size set to $16$ and meticulously document the results in Fig. 5. In our analysis, bubble charts are used to depict the measurement outcomes, wherein the vertical axis denotes the Mean Squared Error (MSE), the horizontal axis quantifies the training duration, and the bubble magnitude correlates with the allocated GPU memory. The visualization reveals that the S-Mamba algorithm attains the most favorable MSE metric across the PEMS07, Electricity, and Traffic datasets. When benchmarked against Transformer-based models, S-Mamba typically necessitates short training time and low allocated GPU memory. The RLinear model does utilize minimal GPU memory and curtails training time, it does not confer a competitive edge in terms of forecast precision. Overall, S-Mamba manifests exemplary predictive accuracy with a low computational resource footprint.
Table 5: Ablation study on Electricity, Traffic, Weather, Solar-Energy, and ETTh2. The lookback length $L=96$, while the forecast length $T\in\left\{96,192,336,720\right\}$.